# Designing LLM Chains by Adapting Techniques from Crowdsourcing Workflows

### v2

@@ 3.3.1. Workflow architecture
Sequential architectures order subtasks in a chain in which the output of one worker passes forward to another worker (Wang et al., 2023b; Nair et al., 2023a; Wu et al., 2022b; Huang et al., 2023a; Zhou et al., 2022b; Huang et al., 2023b; Reppert et al., 2023; Gero et al., 2023; Wu et al., 2023c; Long et al., 2023; Eisenstein et al., 2022).
@@
Branching workflows contain conditionals that, depending on the output of a prior step, choose different execution paths (Huang et al., 2018, 2023b; Kittur et al., 2012; Wu et al., 2022a; Zharikova et al., 2023). Branching architectures can grant flexibility for workflows to address a variety of tasks.
@@
Redundant architectures require different workers to complete the same task in parallel or iteratively. In parallel generation, multiple workers perform the same subtask separately (Weir et al., 2015; Guo et al., 2019; Cobbe et al., 2021; Butler et al., 2017; Lin et al., 2014a; Alshaibani and Quinn, 2021; Huang et al., 2017). This subtask is often, but not always, followed by an aggregation or evaluation subtask to merge or select among the different parallel generations. Iterative architectures run iteratively over a changing input. For example, iterative improvement takes the output of a subtask and iteratively gives it to a series of workers to improve, sometimes with interspersed decision tasks choosing between the current and prior iterations (Luther et al., 2015; André et al., 2014; LaToza et al., 2014; Goto et al., 2016; Dai et al., 2013; Lasecki et al., 2015; Willett et al., 2012). The number of iterations can be determined to be a fixed number of steps (Wu et al., 2023c; Kim et al., 2017; Huffaker et al., 2020; Little et al., 2010b; Zaidan and Callison-Burch, 2011), by including a subtask to check for sufficiency (Madaan et al., 2023b; Creswell et al., 2022; Creswell and Shanahan, 2022; Nair et al., 2023b; Wang et al., 2022; Wu et al., 2023b; Liang et al., 2023; Salehi et al., 2017; Kulkarni et al., 2014; Guo et al., 2016), until the user stops calling iterations (Cheng and Bernstein, 2015; Reppert et al., 2023; Kulkarni et al., 2014; Chilton et al., 2019; Willett et al., 2012), until the outputs converge (Schick et al., 2022; Liem et al., 2011; Du et al., 2023; Gero et al., 2023; Liem et al., 2011), until all quality checks have all been addressed (Kim et al., 2017; Bragg et al., 2013; Chilton et al., 2013; Marcus et al., 2011a, b; Yang et al., 2022; Teevan et al., 2016), from increasing complexity levels (Alshaibani and Quinn, 2021; Alshaibani et al., 2020; Bursztyn et al., 2022), and through decision-theoretic approaches (Park et al., 2022; Dai et al., 2010). Redundant architectures most explicitly support the diverse responses strategy, but they can also be used for quality control (e.g. determining if a disagreement has been resolved by checking similarity across parallel workstreams) and for surfacing uncertainty for user revision.
@@
Dynamic architectures support changing the workflow architecture as it is executed. Map-reduce is a dynamic architecture in which workers break down a task recursively until the subtasks are an appropriate size (Wu et al., 2023c; Liem et al., 2011; Kittur et al., 2011; Kim and Monroy-Hernandez, 2016; Kulkarni et al., 2012; Huang et al., 2023c; Kaur et al., 2017; Valentine et al., 2017; Liem et al., 2011). The literature defines such workflows as a hierarchy (Huang et al., 2023b; Kittur et al., 2011). Other dynamic workflow architectures include doing beam search across multiple options (Xie et al., 2023; Yao et al., 2023; Creswell and Shanahan, 2022) and allowing for workers to flexibly create tasks on a shared to-do list before choosing which of those to complete (Chilton et al., 2019; Nebeling et al., 2016; Mahyar et al., 2018; Zhang et al., 2012; Kittur et al., 2011; LaToza et al., 2014). Some other architecture methods may have dynamic elements, such as parallelizing an input into chunks and running the workflow for each chunk (André et al., 2013; Gouravajhala et al., 2018; Zaidan and Callison-Burch, 2011; Kim et al., 2014; Pavel et al., 2014; Kaur et al., 2017; Bernstein et al., 2010) or dynamically determining the number of iterative steps (Chen et al., 2019; Chilton et al., 2014; Madaan et al., 2023b; Zhou et al., 2022b; Nair et al., 2023b). Dynamic architectures support the workflow flexibility strategy, and can be used to fit within resource constraints, create appropriate subtasks, and incorporate user revision.

### v3

@@ 4.2.3. Workflow architecture
Sequential architectures order subtasks in a chain in which the output of one worker passes forward to another worker (Wang et al., 2023; Nair et al., 2023a; Wu et al., 2022b; Huang et al., 2023a; Zhou et al., 2022b; Huang et al., 2023b; Reppert et al., 2023; Gero et al., 2023; Wu et al., 2023c; Long et al., 2023; Eisenstein et al., 2022).
@@
Branching workflows apply conditionals to choose different execution paths based on the outputs of prior steps (Huang et al., 2018, 2023b; Kittur et al., 2012; Wu et al., 2022a; Zharikova et al., 2023).
@@
Redundant architectures require different workers to complete the same task in parallel or iteratively. In parallel generation, multiple workers perform the same subtask separately (Weir et al., 2015; Guo et al., 2019; Cobbe et al., 2021; Butler et al., 2017; Lin et al., 2014a; Alshaibani and Quinn, 2021; Huang et al., 2017). This subtask is often, but not always, followed by an aggregation or evaluation subtask to merge or select among the different parallel generations. Iterative architectures run iteratively over a changing input. For example, iterative improvement takes the output of a subtask and iteratively gives it to a series of workers to improve, sometimes with interspersed decision tasks choosing between the current and prior iterations (Luther et al., 2015; André et al., 2014; LaToza et al., 2014; Goto et al., 2016; Dai et al., 2013; Lasecki et al., 2015; Willett et al., 2012). The number of iterations can be determined to be a fixed number of steps (Wu et al., 2023c; Kim et al., 2017; Huffaker et al., 2020; Little et al., 2010b; Zaidan and Callison-Burch, 2011), by including a subtask to check for sufficiency (Madaan et al., 2023b; Creswell et al., 2022; Creswell and Shanahan, 2022; Nair et al., 2023b; Wang et al., 2022; Wu et al., 2023b; Liang et al., 2023; Salehi et al., 2017; Kulkarni et al., 2014; Guo et al., 2016), until the user stops calling iterations (Cheng and Bernstein, 2015; Reppert et al., 2023; Kulkarni et al., 2014; Chilton et al., 2019; Willett et al., 2012), until the outputs converge (Schick et al., 2022; Liem et al., 2011; Du et al., 2023; Gero et al., 2023; Liem et al., 2011), until all quality checks are passed (Kim et al., 2017; Bragg et al., 2013; Chilton et al., 2013; Marcus et al., 2011a, b; Yang et al., 2022; Teevan et al., 2016), from increasing complexity levels (Alshaibani and Quinn, 2021; Alshaibani et al., 2020; Bursztyn et al., 2022), and through decision-theoretic approaches (Park et al., 2022; Dai et al., 2010).
@@
Dynamic architectures support changing the workflow architecture as it is executed. Map-reduce is a dynamic architecture in which workers break down a task recursively until the subtasks are an appropriate size (Wu et al., 2023c; Liem et al., 2011; Kittur et al., 2011; Kim and Monroy-Hernandez, 2016; Kulkarni et al., 2012; Huang et al., 2023c; Kaur et al., 2017; Valentine et al., 2017; Liem et al., 2011). The literature defines such workflows as a hierarchy (Huang et al., 2023b; Kittur et al., 2011). Other dynamic workflow architectures include doing beam search across multiple options (Xie et al., 2023; Yao et al., 2023; Creswell and Shanahan, 2022) and allowing for workers to flexibly create tasks on a shared to-do list before choosing which of those to complete (Chilton et al., 2019; Nebeling et al., 2016; Mahyar et al., 2018; Zhang et al., 2012; Kittur et al., 2011; LaToza et al., 2014). Some other architecture methods may have dynamic elements, such as parallelizing an input into chunks and running the workflow for each chunk (André et al., 2013; Gouravajhala et al., 2018; Zaidan and Callison-Burch, 2011; Kim et al., 2014; Pavel et al., 2014; Kaur et al., 2017; Bernstein et al., 2010) or dynamically determining the number of iterative steps (Chen et al., 2019; Chilton et al., 2014; Madaan et al., 2023b; Zhou et al., 2022b; Nair et al., 2023b).